OpenStudy (sleepyjess):

Verify the identity. \(\sf cot(x-\dfrac{\pi}{2})=-tan~x\)

OpenStudy (sleepyjess):

@thomaster

OpenStudy (sleepyjess):

@Directrix

OpenStudy (mathmath333):

use this identities \(\large\tt \begin{align} \color{black}{ \cot x=\dfrac{1}{\tan x} \\~\\ \tan (A+B)=\dfrac{\tan A+\tan B}{1-\tan A\tan B} }\end{align}\)

Directrix (directrix):

I think you need to apply this formula. Maybe @mathmath333 will post it so we can see it better.

Directrix (directrix):

This identity should fall right out with that.

OpenStudy (sleepyjess):

Like this? \(\sf cot(a-b)=\dfrac{1+tan~a~tan~b}{tan~a-tan~b}\)

Directrix (directrix):

Yes. Thanks. I'll draw (I'm low tech) how to begin.

Directrix (directrix):